OpenStudy (anonymous):

Can you show me a example of Parabola, hyperbola, ellipse, and circle equations...

OpenStudy (anonymous):

circle= X^2+Y^2=R^2 parabol= Y=aX^2+bX+c ellipse=(x^2/a^2)+(Y^2/b^2)=1 hypervola=(x^2/a^2)-(Y^2/b^2)=1